# Imputation-free Learning of Tabular Data with Missing Values using Incremental Feature Partitions in Transformer

### v4

@@ 5.1 Robustness to missing value types
The proposed method brings several advantages over existing imputation methods. First, imputation methods are usually not designed to consider and customize for different types of missing values. Therefore, an effective imputation method for MCAR-type data may not be optimal for MNAR-type data or data that naturally come with missing values. The imputation error and variance in the data due to varying missing value types can be avoided using our non-imputation classification method. Moreover, NIIL maximizes the use of observed values in learning, preserving data integrity. This characteristic is particularly valuable in sensitive domains such as healthcare and scientific research, where model-generated synthetic values are often considered untrustworthy and can compromise the credibility of data-driven outcomes. Second, imputation of heterogeneous feature space in tabular data (e.g. numerical, categorical) requires different learning frameworks. For example, continuous numerical value imputation is a regression problem, whereas categorical variables need classification models. The proposed method alleviates the need for separate computational steps and the requirement of train-test splits for the imputation process.
@@ 5.2 Robustness to high missing rates
Imputation-based methods typically learn missing values from their relationship with observed values. However, as the missing value rate increases, the relationship ceases to exist, causing a decline in imputation quality and downstream predictive accuracy. Many existing studies evaluate imputation methods up to a 30% missing rate [42, 43]. The robustness of many similar methods to higher missing rates remains unknown. Real-world scenarios may include important features missing in more than 50% of the samples, as observed in [44]. Imputation methods suffer from higher error rates, resulting in more unreliable data for downstream classification. However, methods without requiring imputation, such as AM-FTT, rely solely on observed values. The proposed NIIL approach avoids the imputation step and instead leverages feature attention via the FTT framework and incremental learning of observed values to remain robust against increasing missing value rates.
@@ 5.3 Comparison with the literature
The literature explores a variety of DL/ML methods for missing-value imputation. A recent survey found that deep generative models (e.g., GAIN, VAE) require a large sample size ($n>30,000$); otherwise, traditional imputation methods (e.g., MICE) tend to outperform generative approaches [18]. Similar requirements for the performance of imputation methods do not apply to NIIL. This may explain why MICE is superior to GAIN in our results for tabular data sets with a standard sample size. Furthermore, the recent diffusion-based method, Diffputer [15] provided a comprehensive benchmarking against multiple advanced baselines, including diffusion-based methods (MissDiff [21], TabCSDI [22]), generative adversarial methods (GAIN [14], MIWAE [45]), iterative imputation (MICE[46], EM, HyperImpute [19]), and graph neural network-based approaches (MIRACLE [47], GRAPE [48]). Diffputer demonstrates better performance than these baseline methods, particularly highlighting its suitability for large datasets ($n>10,000$). Our results suggest that Diffputer performance may not generalize well across different data sets, missing value rates, and types, unlike the proposed NIIL method. It may be inferred that NIIL would also outperform those advanced baseline deep learning methods for imputing missing values. Despite the advantages of deep imputation methods in large-scale datasets, practitioners must carefully consider the missing value rates, types, and data characteristics when selecting appropriate imputation methods for real-world applications.
@@ 5.4 Broader impact
This paper introduces a novel machine learning method for tabular data with missing values. Missing values appear in varying rates and types in many application domains, where one imputation method may not be optimal for all scenarios. Missing values in tabular data sets critically challenge the feasibility and performance of machine learning models. The widespread adoption of arbitrary missing value imputation strategies raises many questions about data quality and information reliability among data stakeholders. Our proposed ML method enables machine learning without missing value imputation, alleviating data wastage, biases, errors, and computational costs associated with missing value imputations. The proposed method foresees positive social impacts through increased inclusivity, integrity, trustworthiness, and efficiency in data analysis workflows. The impact on health science and the analysis of health records data is significant, where missing values are a major barrier to the development of data-driven precision medicine. However, practitioners should consider ethical implications, such as the risks of reinforcing biases in the original data set or overlooking the reasons behind missing data.
@@ 5.5 Limitations
The proposed NIIL method performs remarkably in deep learning tabular data without imputation or initializing missing values. Although deep learning without imputation of missing values is a significant contribution, the proposed method has several limitations. When the missing value rate is low and completely random, sophisticated methods for imputing missing values can yield better results than our proposed non-imputation method. Several imputation and classification models are computationally more efficient than the proposed non-imputation classification method. However, the difference in runtime is around several minutes for tabular data sets of standard sizes. When data quality and trustworthiness matter, especially in the most practical missing value scenarios (naturally occurring and MNAR type), the proposed NIIL method results in the overall best performance.
6 Conclusions
This paper introduces NIIL, a novel incremental learning algorithm that is performance-wise effective, robust to missing value rates and types, and computationally efficient for learning tabular data with missing values. Although attention masks in the feature-tokenized transformer enable no imputation learning of tabular data with missing values, incremental learning additionally yields substantial and consistent performance improvement. The proposed NIIL algorithm outperforms state-of-the-art deep imputation methods in downstream classification tasks, in addition to its ability to learn the high-dimensional feature space incrementally. Empirically, we show that the most computationally efficient and performance-wise effective feature partition size for incremental learning is half of the feature dimension of the data set.
@@ 7 Acknowledgments
Appendix A Number of operations and computational cost
@@
Appendix B Data set-specific AUC classification scores
@@
Dataset ID % Median-GBT Median-FTT MICE-GBT MICE-FTT GAIN-GBT GAIN-FTT Diffputer-FTT Diffputer-GBT MIA-Xgboost MIA-LightGBM AM-FTT NIIL (K=d/2)
1063 10 0.840 (0.06) 0.840 (0.05) 0.837 (0.04) 0.838 (0.04) 0.795 (0.04) 0.846 (0.04) 0.847 (0.04) 0.820 (0.05) 0.791 (0.05) 0.795 (0.07) 0.856 (0.04) 0.862 (0.04)
50 0.802 (0.05) 0.732 (0.02) 0.824 (0.06) 0.833 (0.04) 0.772 (0.06) 0.836 (0.05) 0.849 (0.05) 0.785 (0.08) 0.780 (0.05) 0.773 (0.05) 0.860 (0.04) 0.854 (0.04)
463 10 0.642 (0.16) 0.717 (0.21) 0.653 (0.06) 0.694 (0.09) 0.672 (0.11) 0.617 (0.24) 0.566 (0.21) 0.654 (0.12) 0.634 (0.08) 0.630 (0.06) 0.587 (0.25) 0.728 (0.05)
50 0.664 (0.09) 0.621 (0.21) 0.565 (0.09) 0.714 (0.11) 0.601 (0.03) 0.596 (0.27) 0.481 (0.18) 0.593 (0.13) 0.583 (0.13) 0.501 (0.11) 0.523 (0.14) 0.674 (0.10)
13 10 0.713 (0.07) 0.505 (0.10) 0.713 (0.07) 0.691 (0.05) 0.659 (0.07) 0.704 (0.06) 0.647 (0.07) 0.670 (0.07) 0.628 (0.08) 0.643 (0.04) 0.661 (0.07) 0.660 (0.07)
50 0.672 (0.09) 0.489 (0.08) 0.672 (0.09) 0.623 (0.04) 0.628 (0.05) 0.652 (0.07) 0.561 (0.08) 0.628 (0.05) 0.588 (0.10) 0.629 (0.06) 0.488 (0.10) 0.571 (0.11)
31 10 0.761 (0.03) 0.747 (0.03) 0.749 (0.04) 0.756 (0.03) 0.737 (0.03) 0.733 (0.02) 0.761 (0.04) 0.743 (0.02) 0.743 (0.03) 0.680 (0.04) 0.746 (0.03) 0.753 (0.03)
50 0.657 (0.03) 0.658 (0.06) 0.646 (0.04) 0.704 (0.04) 0.705 (0.04) 0.693 (0.05) 0.695 (0.03) 0.657 (0.03) 0.666 (0.05) 0.642 (0.03) 0.656 (0.05) 0.674 (0.04)
37 10 0.836 (0.03) 0.841 (0.04) 0.841 (0.03) 0.772 (0.04) 0.794 (0.02) 0.803 (0.03) 0.804 (0.03) 0.805 (0.03) 0.812 (0.03) 0.823 (0.03) 0.804 (0.03) 0.804 (0.03)
50 0.754 (0.05) 0.763 (0.06) 0.724 (0.05) 0.687 (0.05) 0.737 (0.02) 0.548 (0.10) 0.683 (0.06) 0.696 (0.07) 0.706 (0.04) 0.716 (0.04) 0.703 (0.04) 0.718 (0.05)
23381 10 0.577 (0.05) 0.580 (0.03) 0.588 (0.05) 0.582 (0.03) 0.621 (0.04) 0.576 (0.09) 0.593 (0.06) 0.570 (0.05) 0.530 (0.07) 0.557 (0.05) 0.603 (0.02) 0.623 (0.03)
50 0.523 (0.06) 0.488 (0.10) 0.520 (0.06) 0.488 (0.10) 0.527 (0.05) 0.561 (0.04) 0.557 (0.05) 0.560 (0.09) 0.509 (0.03) 0.569 (0.04) 0.476 (0.06) 0.516 (0.03)
1480 10 0.727 (0.03) 0.708 (0.04) 0.734 (0.05) 0.742 (0.04) 0.731 (0.02) 0.724 (0.05) 0.723 (0.07) 0.733 (0.03) 0.720 (0.05) 0.738 (0.05) 0.663 (0.04) 0.729 (0.05)
50 0.715 (0.07) 0.620 (0.09) 0.684 (0.07) 0.696 (0.02) 0.679 (0.03) 0.650 (0.05) 0.641 (0.06) 0.680 (0.03) 0.665 (0.04) 0.669 (0.05) 0.596 (0.06) 0.678 (0.06)
1067 10 0.800 (0.02) 0.791 (0.02) 0.810 (0.02) 0.793 (0.02) 0.799 (0.02) 0.796 (0.02) 0.807 (0.02) 0.806 (0.03) 0.992 (0.01) 0.993 (0.00) 0.797 (0.02) 0.875 (0.16)
50 0.778 (0.02) 0.695 (0.03) 0.792 (0.03) 0.781 (0.02) 0.784 (0.03) 0.788 (0.01) 0.785 (0.02) 0.771 (0.03) 0.796 (0.04) 0.801 (0.03) 0.724 (0.02) 0.793 (0.02)
334 10 0.684 (0.07) 0.721 (0.13) 0.684 (0.07) 0.805 (0.12) 0.628 (0.09) 0.856 (0.01) 0.721 (0.13) 0.717 (0.08) 0.803 (0.04) 0.756 (0.06) 0.822 (0.11) 0.853 (0.08)
50 0.528 (0.06) 0.531 (0.04) 0.528 (0.06) 0.565 (0.04) 0.537 (0.09) 0.577 (0.07) 0.531 (0.04) 0.532 (0.04) 0.603 (0.01) 0.620 (0.04) 0.553 (0.02) 0.598 (0.07)
1071 10 0.650 (0.10) 0.682 (0.19) 0.778 (0.07) 0.778 (0.08) 0.657 (0.11) 0.795 (0.06) 0.734 (0.10) 0.710 (0.12) 0.762 (0.05) 0.760 (0.03) 0.712 (0.15) 0.798 (0.03)
50 0.695 (0.06) 0.665 (0.21) 0.690 (0.08) 0.746 (0.11) 0.685 (0.07) 0.767 (0.09) 0.741 (0.09) 0.650 (0.07) 0.650 (0.09) 0.676 (0.08) 0.667 (0.15) 0.742 (0.06)
1498 10 0.700 (0.03) 0.755 (0.03) 0.701 (0.03) 0.744 (0.04) 0.698 (0.06) 0.745 (0.05) 0.738 (0.03) 0.704 (0.05) 0.671 (0.04) 0.675 (0.03) 0.733 (0.04) 0.738 (0.04)
50 0.656 (0.03) 0.693 (0.03) 0.668 (0.05) 0.731 (0.04) 0.683 (0.03) 0.707 (0.03) 0.649 (0.06) 0.684 (0.06) 0.631 (0.06) 0.641 (0.05) 0.654 (0.09) 0.652 (0.08)
40691 10 0.770 (0.03) 0.746 (0.01) 0.789 (0.03) 0.704 (0.04) 0.758 (0.05) 0.735 (0.02) 0.733 (0.02) 0.736 (0.05) 0.834 (0.02) 0.829 (0.02) 0.724 (0.02) 0.744 (0.02)
50 0.692 (0.04) 0.669 (0.01) 0.685 (0.04) 0.644 (0.02) 0.698 (0.01) 0.682 (0.02) 0.604 (0.02) 0.604 (0.05) 0.699 (0.03) 0.718 (0.04) 0.654 (0.02) 0.664 (0.01)
181 10 0.819 (0.04) 0.792 (0.02) 0.798 (0.05) 0.744 (0.02) 0.803 (0.03) 0.808 (0.02) 0.784 (0.02) 0.803 (0.04) 0.832 (0.03) 0.835 (0.03) 0.738 (0.01) 0.768 (0.01)
50 0.686 (0.06) 0.716 (0.02) 0.701 (0.06) 0.642 (0.02) 0.710 (0.05) 0.505 (0.02) 0.634 (0.02) 0.698 (0.06) 0.731 (0.05) 0.724 (0.05) 0.642 (0.02) 0.704 (0.01)

### v5

@@ 5.1 Robustness to missing value types
The proposed method brings several advantages over existing imputation methods. First, imputation methods are usually not selected and designed based on different missing value types. Therefore, an effective imputation method for MCAR-type data may not be optimal for MNAR-type data or data that have natural missing values. The imputation error and variance in the data due to varying missing value types can be avoided using our imputation-free learning method. Furthermore, IFIAL maximizes the use of observed values in learning, thereby preserving data integrity. This characteristic is particularly valuable in sensitive domains such as healthcare and scientific research, where model-generated synthetic values are often considered unreliable and can compromise the credibility of data-driven outcomes. Second, imputation of heterogeneous feature space in tabular data (e.g. numerical, categorical) requires different learning frameworks. For example, continuous numerical value imputation is a regression problem, whereas categorical variables need classification models. The proposed method eliminates the need for separate computational steps and the requirement of train-test splits to build and evaluate the imputation model.
@@ 5.2 Robustness to high missing rates
Imputation-based methods typically learn missing values from their relationship with observed values. However, as the missing value rate increases, the relationship ceases to exist, causing a decline in imputation quality and downstream predictive accuracy. Many existing studies evaluate imputation methods up to a 30% missing rate [lee2023, aimnet]. The robustness of many similar methods to higher missing rates remains unknown. Real-world scenarios may include important features missing in more than 50% of the samples, as observed in [Samad2019Jaac]. Imputation methods are prone to higher error rates, leading to more unreliable data for downstream classification. However, methods without requiring imputation, such as AM-FTT, rely solely on observed values. The proposed IFIAL approach skips the imputation step and instead leverages feature attention via the FTT framework and incremental learning of observed values to remain robust against increasing missing value rates.
@@ 5.3 Comparison with the literature
The literature explores a variety of DL/ML methods for missing-value imputation. A recent survey found that deep generative models (e.g., GAIN, VAE) require a large sample size ($n>30,000$); otherwise, traditional imputation methods (e.g., MICE) tend to outperform generative approaches [sun2023deep]. Similar sample size requirements for imputation methods do not apply to IFIAL. This may explain why MICE is superior to GAIN in our results for tabular data sets with a decent sample size. Furthermore, the recent diffusion-based method, Diffputer [diffputer] provides a comprehensive benchmarking against multiple advanced baselines, including diffusion-based methods (MissDiff [ouyang2023missdiff], TabCSDI [tabcsdi]), generative adversarial methods (GAIN [yoon2018gain], MIWAE [mattei2019miwae]), iterative imputation (MICE[mice], EM, HyperImpute [hyperimpute]), and graph neural network-based approaches (MIRACLE [kyono2021miracle], GRAPE [grape]). Diffputer demonstrates better performance than these baselines, particularly highlighting its suitability for large data sets ($n>10,000$). Our results suggest that Diffputer performance may not generalize well across different data sets, missing value rates, and types, unlike the proposed IFIAL method. It may be inferred that IFIAL would also outperform advanced imputation baselines used in benchmarking the Diffputer performance. Despite the advantages of deep imputation methods in large-scale datasets, practitioners must carefully consider the missing value rates, types, and data characteristics when selecting appropriate imputation methods for real-world applications. In particular, the choice of the imputation technique affects the quality of the data and can produce varying classification results for the same dataset and classifier pair. For example, in data set ID 463 with 50% missingness of MCAR type, the FTT AUC score decreases markedly from 0.670 to 0.571 when median imputation is applied instead of MICE. In contrast, for data set ID 31 with 10% missingness of MNAR type, the AUC score drops from 0.745 to 0.687 when MICE imputation is used rather than the median. The complete removal of an imputation step, as suggested by IFIAL, preserves data quality while still delivering strong and competitive classification performance.
@@ 5.4 Broader impact
This paper introduces a novel machine learning method for tabular data with missing values. Missing values appear in varying rates and types in many application domains, where one imputation method may not be optimal for all scenarios. Missing values in tabular data sets critically challenge the feasibility and performance of machine learning models. The widespread adoption of arbitrary missing value imputation strategies raises many questions about data quality and information reliability among data stakeholders. Our proposed ML method enables machine learning without missing value imputation, alleviating data wastage, biases, errors, and computational costs associated with missing value imputations. The proposed method is expected to have positive social impacts through greater inclusivity, integrity, trustworthiness, and efficiency in data analysis workflows. The impact on health science and the analysis of health records data is significant, where missing values are a major barrier to the development of data-driven precision medicine. However, practitioners should consider ethical implications, such as the risks of reinforcing biases in the original data set or overlooking the reasons behind missing data.
@@ 5.5 Limitations
The proposed IFIAL method performs remarkably well in deep learning with tabular data, eliminating the need to impute and initialize missing values. Although deep learning without imputation of missing values is a significant contribution, the proposed method has several limitations. When the missing value rate is low and completely random, sophisticated methods for imputing missing values can yield better results than our proposed imputation-free method. Several imputation and classification models are computationally more efficient than the proposed non-imputation classification method. The performance gain of the proposed method over other methods diminishes when the sample size is very large, particularly at a low missing rate.
6 Conclusion
This paper introduces IFIAL, a novel imputation-free deep learning algorithm that is performance-wise effective, robust to missing value rates and types, and computationally efficient for learning tabular data with missing values. Although attention masks in the feature tokenized transformer enable imputation-free learning of tabular data with missing values, incremental learning in feature partitions additionally yields better performance and computational efficiency. The proposed IFIAL algorithm outperforms state-of-the-art deep imputation methods in downstream classification tasks, in addition to its ability to learn the high-dimensional feature space in small feature partitions. Empirically, we demonstrate that the most computationally efficient and performance-wise adequate feature partition size for incremental learning is half the dimension of the feature space of the data set. Future work should aim at investigating imputation-free solutions for other deep learning methods.
@@ 7 Acknowledgments
Appendix A Optuna Configurations
Table 6: Hyperparameter search space used in Optuna for IFIAL, FTT and tree-based baselines.
Parameter Distribution
IFIAL and FTT
   num_attention_head Categorical{4, 8}
   hidden_dim Categorical{64, 128, 256}
   num_layer UniformInt[2, 6]
   ffn_ratio Categorical{2, 4}
   activation Categorical{”relu”, ”gelu”, ”leakyrelu”}
   hidden_dropout_prob UniformFloat[0.0, 0.5] (step=0.1)
   weight_decay LogUniform[1e-6, 1e-2]
   learning_rate LogUniform[1e-5, 1e-3]
   batch_size Categorical{32, 64, 128, 256}
   warmup_ratio Categorical{0.0, 0.05, 0.1}
Gradient Boosting Trees (GBT)
   n_estimators UniformInt[50, 10000]
   learning_rate LogUniform[1e-3, 0.3]
   max_depth UniformInt[1, 10]
   min_samples_split UniformInt[2, 20]
   min_samples_leaf UniformInt[1, 20]
   subsample UniformFloat[0.2, 1.0]
   max_features UniformFloat[0.2, 1.0]
XGBoost
   max_depth UniformInt[1, 10]
   learning_rate LogUniform[$\exp(-7)$, 1]
   n_estimators UniformInt[100, 4000]
   subsample UniformFloat[0.2, 1.0]
   colsample_bytree UniformFloat[0.2, 1.0]
   min_child_weight LogUniform[$\exp(-16)$, $\exp(5)$]
   gamma LogUniform[$\exp(-16)$, $\exp(2)$]
   reg_alpha LogUniform[$\exp(-16)$, $\exp(2)$]
   reg_lambda LogUniform[$\exp(-16)$, $\exp(2)$]
LightGBM
   n_estimators UniformInt[50, 10000]
   learning_rate LogUniform[1e-3, 0.3]
   num_leaves UniformInt[20, 200]
   max_depth UniformInt[-1, 15]
   min_child_samples UniformInt[5, 100]
   subsample UniformFloat[0.2, 1.0]
   subsample_freq UniformInt[0, 10]
   colsample_bytree UniformFloat[0.2, 1.0]
   reg_alpha LogUniform[1e-6, 10]
   reg_lambda LogUniform[1e-6, 10]
   min_split_gain LogUniform[1e-6, 1]
Appendix B Number of operations and computational cost
@@
Appendix C Data set-specific AUC classification scores
@@
Dataset ID % Median-GBT Median-FTT MICE-GBT MICE-FTT GAIN-GBT GAIN-FTT Diffputer-FTT Diffputer-GBT MIA-Xgboost MIA-LightGBM AM-FTT IFIAL (K=d/2)
13 10 0.653 (0.08) 0.639 (0.09) 0.661 (0.06) 0.641 (0.05) 0.675 (0.08) 0.654 (0.07) 0.597 (0.07) 0.658 (0.08) 0.646 (0.12) 0.645 (0.08) 0.611 (0.08) 0.598 (0.17)
50 0.669 (0.08) 0.555 (0.11) 0.648 (0.14) 0.573 (0.05) 0.643 (0.10) 0.602 (0.08) 0.511 (0.09) 0.650 (0.08) 0.626 (0.12) 0.600 (0.13) 0.538 (0.12) 0.580 (0.14)
31 10 0.705 (0.02) 0.697 (0.03) 0.714 (0.03) 0.706 (0.04) 0.707 (0.02) 0.683 (0.02) 0.711 (0.04) 0.690 (0.02) 0.718 (0.03) 0.637 (0.02) 0.696 (0.03) 0.736 (0.04)
50 0.698 (0.03) 0.608 (0.06) 0.689 (0.04) 0.654 (0.04) 0.695 (0.04) 0.643 (0.05) 0.645 (0.03) 0.675 (0.02) 0.712 (0.03) 0.596 (0.03) 0.606 (0.05) 0.655 (0.05)
37 10 0.586 (0.03) 0.791 (0.04) 0.589 (0.03) 0.722 (0.04) 0.624 (0.04) 0.753 (0.04) 0.754 (0.03) 0.556 (0.05) 0.654 (0.02) 0.558 (0.04) 0.754 (0.04) 0.630 (0.10)
50 0.561 (0.05) 0.713 (0.06) 0.593 (0.04) 0.637 (0.06) 0.584 (0.02) 0.598 (0.11) 0.633 (0.07) 0.686 (0.06) 0.675 (0.05) 0.676 (0.07) 0.653 (0.04) 0.555 (0.08)
181 10 0.601 (0.03) 0.742 (0.02) 0.645 (0.05) 0.694 (0.02) 0.665 (0.05) 0.758 (0.02) 0.734 (0.03) 0.617 (0.03) 0.663 (0.04) 0.649 (0.04) 0.688 (0.01) 0.743 (0.01)
50 0.560 (0.02) 0.666 (0.03) 0.590 (0.02) 0.592 (0.02) 0.598 (0.04) 0.655 (0.02) 0.584 (0.02) 0.587 (0.03) 0.556 (0.06) 0.549 (0.00) 0.592 (0.02) 0.722 (0.02)
334 10 0.647 (0.08) 0.671 (0.15) 0.673 (0.05) 0.755 (0.13) 0.666 (0.08) 0.906 (0.02) 0.671 (0.15) 0.643 (0.07) 0.670 (0.05) 0.643 (0.07) 0.772 (0.12) 0.912 (0.01)
50 0.699 (0.09) 0.681 (0.04) 0.670 (0.04) 0.515 (0.05) 0.665 (0.07) 0.527 (0.07) 0.681 (0.04) 0.637 (0.05) 0.674 (0.08) 0.637 (0.05) 0.503 (0.03) 0.620 (0.08)
463 10 0.526 (0.19) 0.667 (0.04) 0.679 (0.19) 0.644 (0.10) 0.608 (0.09) 0.546 (0.31) 0.516 (0.24) 0.577 (0.20) 0.655 (0.12) 0.516 (0.24) 0.673 (0.18) 0.747 (0.08)
50 0.678 (0.15) 0.571 (0.04) 0.670 (0.09) 0.664 (0.12) 0.667 (0.14) 0.567 (0.27) 0.631 (0.20) 0.552 (0.16) 0.605 (0.12) 0.552 (0.16) 0.537 (0.30) 0.666 (0.10)
1063 10 0.699 (0.10) 0.790 (0.06) 0.507 (0.05) 0.788 (0.05) 0.675 (0.09) 0.796 (0.05) 0.797 (0.05) 0.697 (0.04) 0.686 (0.03) 0.507 (0.05) 0.806 (0.05) 0.771 (0.11)
50 0.793 (0.08) 0.682 (0.02) 0.807 (0.04) 0.783 (0.05) 0.790 (0.04) 0.786 (0.05) 0.799 (0.05) 0.821 (0.05) 0.845 (0.03) 0.682 (0.02) 0.810 (0.05) 0.805 (0.05)
1067 10 0.769 (0.02) 0.741 (0.02) 0.936 (0.01) 0.743 (0.02) 0.897 (0.02) 0.746 (0.02) 0.757 (0.02) 0.937 (0.01) 0.936 (0.01) 0.741 (0.02) 0.747 (0.03) 0.808 (0.03)
50 0.699 (0.03) 0.645 (0.04) 0.750 (0.04) 0.731 (0.02) 0.736 (0.03) 0.738 (0.01) 0.735 (0.02) 0.757 (0.03) 0.699 (0.03) 0.645 (0.04) 0.674 (0.02) 0.794 (0.02)
1071 10 0.651 (0.09) 0.632 (0.22) 0.634 (0.13) 0.728 (0.09) 0.662 (0.16) 0.745 (0.07) 0.684 (0.11) 0.651 (0.09) 0.634 (0.13) 0.553 (0.11) 0.662 (0.16) 0.755 (0.09)
50 0.669 (0.10) 0.615 (0.23) 0.663 (0.13) 0.696 (0.12) 0.617 (0.17) 0.717 (0.10) 0.691 (0.10) 0.638 (0.14) 0.617 (0.17) 0.607 (0.14) 0.617 (0.17) 0.648 (0.10)
1480 10 0.662 (0.05) 0.658 (0.05) 0.626 (0.03) 0.692 (0.05) 0.658 (0.05) 0.674 (0.06) 0.673 (0.08) 0.655 (0.07) 0.626 (0.03) 0.613 (0.04) 0.613 (0.04) 0.707 (0.07)
50 0.503 (0.09) 0.570 (0.10) 0.652 (0.05) 0.646 (0.02) 0.546 (0.06) 0.600 (0.06) 0.591 (0.06) 0.627 (0.05) 0.546 (0.06) 0.503 (0.09) 0.546 (0.06) 0.715 (0.09)
1498 10 0.642 (0.06) 0.705 (0.03) 0.655 (0.06) 0.694 (0.05) 0.683 (0.05) 0.695 (0.06) 0.688 (0.03) 0.639 (0.05) 0.642 (0.06) 0.639 (0.07) 0.683 (0.05) 0.696 (0.07)
50 0.646 (0.07) 0.643 (0.03) 0.670 (0.04) 0.681 (0.04) 0.604 (0.10) 0.657 (0.03) 0.599 (0.06) 0.472 (0.07) 0.599 (0.06) 0.546 (0.06) 0.604 (0.10) 0.678 (0.03)
23381 10 0.551 (0.06) 0.532 (0.03) 0.564 (0.02) 0.532 (0.03) 0.543 (0.07) 0.526 (0.10) 0.543 (0.07) 0.572 (0.05) 0.532 (0.03) 0.526 (0.10) 0.553 (0.02) 0.612 (0.03)
50 0.560 (0.06) 0.638 (0.11) 0.565 (0.05) 0.638 (0.11) 0.564 (0.03) 0.511 (0.05) 0.507 (0.05) 0.564 (0.03) 0.511 (0.05) 0.507 (0.05) 0.626 (0.07) 0.555 (0.05)
40691 10 0.627 (0.01) 0.696 (0.01) 0.613 (0.04) 0.654 (0.05) 0.683 (0.02) 0.685 (0.02) 0.683 (0.02) 0.582 (0.03) 0.683 (0.02) 0.654 (0.05) 0.674 (0.03) 0.722 (0.01)
50 0.594 (0.05) 0.619 (0.01) 0.568 (0.02) 0.594 (0.02) 0.604 (0.02) 0.632 (0.02) 0.554 (0.03) 0.562 (0.04) 0.594 (0.02) 0.554 (0.03) 0.604 (0.02) 0.687 (0.02)